OpenStudy (soahelp):

Am i right that the answer is C If 2/5 of a ton of concrete covers 5/6 of a wall, how many tons of concrete are required to cover the entire wall? 12/25 ton 3/5 ton 1 and 7/30 tons 2 and 1/12 tons

OpenStudy (igreen):

No..that is incorrect.

OpenStudy (soahelp):

how

OpenStudy (soahelp):

it would be a i see what i did wround

OpenStudy (igreen):

Just divide them. \(\sf \dfrac{2}{5} \div \dfrac{5}{6}\) FLip the 2nd fraction and multiply: \(\sf \dfrac{2}{5} \times \dfrac{6}{5}\) Can you multiply?

OpenStudy (igreen):

Yep, A is correct.. \(\sf \dfrac{2}{5} \times \dfrac{6}{5} \rightarrow \dfrac{2\times 6}{5\times 5} \rightarrow \dfrac{12}{25}\)

OpenStudy (soahelp):

thxs for your help

OpenStudy (phi):

one way to figure out what to do is think "ratio" 2/5 ton "is to" 5/6 wall "as" x tons "is to" 1 wall (notice 2/5 "goes with" 5/6 of a wall, and we want x (for unknown) "goes with" 1 full wall write the ratio as fractions: \[ \frac{ \frac{2}{5}}{\frac{5}{6}} = \frac{x}{1} \] of course x/1 is just x so the problem is \[ x= \frac{ \frac{2}{5}}{\frac{5}{6}} \]

OpenStudy (phi):

to simplify the messy fraction, use the idea that you can multiply top and bottom by the same number. in this case , use 6/5, like this \[ x= \frac{ \frac{2}{5}\cdot \frac{6}{5}}{\frac{5}{6}\cdot \frac{6}{5}} \] notice that 5/6 * 6/5 is 1, so you have \[ x=\frac{ \frac{2}{5}\cdot \frac{6}{5}}{1} \\ x= \frac{2}{5}\cdot \frac{6}{5} \]

OpenStudy (phi):

some people remember that you can change divide to multiplying by first "flipping" the bottom number. so the fast way is to remember that \[ x= \frac{ \frac{2}{5}}{\frac{5}{6}} = \frac{2}{5} \cdot \frac{6}{5} \]